# DocKylin: A Large Multimodal Model for Visual Document Understanding with Efficient Visual Slimming

### v1

@@ Abstract
Current multimodal large language models (MLLMs) face significant challenges in visual document understanding (VDU) tasks due to the high resolution, dense text, and complex layouts typical of document images. These characteristics demand a high level of detail perception ability from MLLMs. While increasing input resolution improves detail perception, it also leads to longer sequences of visual tokens, increasing computational costs and straining the models’ ability to handle long contexts. To address these challenges, we introduce DocKylin, a document-centric MLLM that performs visual content slimming at both the pixel and token levels, thereby reducing token sequence length in VDU scenarios. DocKylin utilizes an Adaptive Pixel Slimming (APS) preprocessing module to perform pixel-level slimming, increasing the proportion of informative pixels. Moreover, DocKylin incorporates a novel Dynamic Token Slimming (DTS) module to conduct token-level slimming, filtering essential tokens and removing others to create a compressed, adaptive visual sequence. Experiments demonstrate DocKylin’s promising performance across various VDU benchmarks. Notably, both the proposed APS and DTS are parameter-free, facilitating easy integration into existing MLLMs, and our experiments indicate their potential for broader applications.
1 Introduction
@@
With the substantial growth in data and model sizes, as well as alignment with human preferences, current Large Language Models (LLMs) [45, 65, 63, 2, 62, 73] have demonstrated remarkable capabilities in reasoning, common sense understanding, expertise in specialized knowledge, zero/few-shot learning, and instruction following, shedding light on the potential for Artificial General Intelligence. Multimodal Large Language Models (MLLMs) aim to enhance these powerful LLMs by endowing them with multimodal capabilities, enabling the integration of diverse modalities beyond text alone.
@@
Although there has been significant progress in the development of MLLMs, their performance in visual document understanding (VDU) tasks remains suboptimal [36, 78]. A primary reason for this is their support for input resolutions of only up to 448x448, which is inadequate for high-resolution, fine-grained document images. Numerous efforts [76, 37, 29, 70, 3, 74, 32, 72] have recently been made to increase the input resolution to address this limitation and achieved notable improvements. However, these efforts still face several challenges: 1) Increasing the resolution also escalates the redundant regions within documents, which results in an increased number of redundant visual tokens, complicating the task for LLMs to identify the correct answers. 2) Some methods [29, 70, 72, 3] employ a fixed rectangular resolution, which can lead to text distortion given the typically large aspect ratios of document images. Moreover, scaling smaller images to larger ones results in inefficiencies. 3) To manage the extended sequences of visual tokens resulting from high resolution, most methods [29, 76, 37, 3] employ a fixed compression ratio or extract a predetermined number of tokens. While this approach may be suitable for natural scene images, it is less effective for document scenarios where content density varies significantly. For example, images of scientific papers are information-dense, whereas images of identification cards are comparatively sparse.
@@
To tackle these challenges, we propose a series of solutions. Firstly, we propose a parameter-free Adaptive Pixel Slimming (APS) preprocessing technique that utilizes gradient information to identify and eliminate redundant regions within document images, thereby reducing the proportion of redundant pixels and enhancing computational efficiency. Secondly, we adopt a more flexible image encoding strategy by capping the maximum number of pixels instead of fixing resolutions. This allows for variable resolutions and aspect ratios. Our lightweight Swin encoder [38] processes high-resolution document images holistically, ensuring feature coherence. Thirdly, we propose a Dynamic Token Slimming (DTS) method based on k-means clustering [39], which is also parameter-free. DTS filters useful tokens from a vast pool of visual tokens, obtaining visual sequences with reduced and dynamic lengths. Incorporating these innovations, we developed our DocKylin, which demonstrates superior performance across multiple VDU benchmarks. Ablation studies further confirm the effectiveness of some of these designs and highlight their potential applicability to other MLLMs or scenarios.
2 Related Works
2.1 Text-centric MLLMs
@@
Text is crucial for understanding images, but the ability of existing MLLMs in this regard is generally subpar [36, 51, 61]. Therefore, there has been a surge of recent work studying how to enhance the perception, recognition, and understanding abilities of text images in MLLMs. Some approaches [58, 23] leverage additional Optical Character Recognition (OCR) engines to recognize text in images and input it into MLLMs. Others [29, 37, 3, 12, 13, 74, 70] enhance the fine-grained perceptual capabilities of MLLMs by maintaining the high resolution of the input images. This includes strategies such as splitting the image into small pieces to accommodate the VIT’s requirement for fixed and small-size inputs and utilizing visual encoders that are more suitable for text images. Authors of [14, 69] explore learning text grounding for MLLMs, while [16, 60] investigate the use of more superior and larger scale instruction data to boost performance. Although these efforts have significantly improved the performance of MLLMs on text-centric tasks, there is still substantial room for improvement.
2.2 Visual Token Compression in MLLMs
@@
Maintaining high-resolution input images is a straightforward and effective method to enhance the performance of MLLMs on text images. However, increased resolutions inevitably lead to a longer sequence of visual tokens. There are various methods for compressing the length of visual sequences in current MLLMs. One of the earlier and commonly used approaches [11, 27, 1] involves utilizing a group of learnable query tokens to extract information in a cross-attention manner. Another method [7, 76] involves concatenating adjacent tokens along the channel dimension to reduce sequence length. Additionally, convolutional neural networks (CNNs) have also been proven effective for visual token compression [6, 16]. However, all these methods either extract tokens of a fixed length or compress them at a fixed ratio. This approach is unsuited for document images, where the content density can vary significantly. A token sequence that is too short or a compression ratio that is too high can result in an inadequate representation of some document images. Conversely, longer token sequences or lower compression rates can reduce computational efficiency.
@@
Recent studies [5] have shown that the optimal length of visual token sequences varies for each sample. Consequently, some recent approaches have explored the adaptive compression of visual sequences. Most of these methods [34, 49, 8, 31] dynamically discard nonessential tokens using attention scores between visual tokens or between visual tokens and class tokens. However, considering the differences between text and natural objects with complete semantics, this paradigm may not be suitable for text images. For text-centric images, HRVDA[72] employs text position annotations to train a separate visual token filtering network, which requires introducing additional parameters training procedures. Compared to the methods above, we explore alternative dynamic compression schemes specifically for document images, which include pixel-level redundancy removal and unsupervised clustering for filtering out nonessential tokens.
3 Methodology
@@
As illustrated in Fig. 1, similar to many multimodal large language models (MLLMs), our architecture also incorporates an image encoder, an MLP layer that aligns the visual modality to the language modality, and a LLM. Building on this framework, DocKylin includes the following enhancements: Images are preprocessed by the Adaptive Pixel Slimming module before being fed to the image encoder, which removes redundant regions and reduces resolution while increasing the proportion of informative pixels. The visual encoder takes the resulting image as input and maintains the aspect ratio. The visual tokens obtained by the visual encoder and MLP layer are handled by our Dynamic Token Slimming (DTS) module to remove nonessential tokens to enhance both performance and efficiency. We provide a detailed description of these improvements in the following.
3.1 Adaptive Pixel Slimming
@@
Perceiving textual information in images typically requires maintaining a high resolution of input, especially for text-rich document images. This poses significant challenges for current resource-intensive MLLMs. To address this issue, existing methods employ strategies such as dividing the images into patches for processing [37, 74], employing visual encoders that support high resolution [70], leveraging OCR results as additional inputs [58, 23], or applying compression transformations to the images beforehand [13].
@@
Here, we propose a novel approach called Adaptive Pixel Slimming (APS) to efficiently reduce the resolution of document images, which is orthogonal to the approaches mentioned above. APS is based on the observation that, as illustrated in Fig. 2, while document images usually exhibit high resolution, they commonly contain a significant proportion of redundant regions. These regions not only increase the computational burden on the visual encoder but may also pose challenges [37, 72] to the reasoning capabilities of the LLMs by the increased redundant visual tokens. APS is thus designed to identify and remove these regions. Specifically, as shown in Fig. 2, during the gradient extraction stage, we utilize the Sobel operator to extract the gradient map from the original image . Regions with low gradient values are considered smooth and devoid of significant information. Therefore, in the redundant region determination stage, we identify the redundant regions in the horizontal and vertical directions by locating areas with consistently low gradient values. For instance, to identify redundant regions in the image’s horizontal direction: if the sum of gradient values for a specific row is below a predefined threshold, we classify that row as redundant. A sequence of contiguous redundant rows constitutes a redundant region in the horizontal direction. The computation for the vertical direction follows a similar procedure. The final image can be obtained by removing the identified redundant rows and columns. For more details on the algorithm and parameter settings, please refer to the Appendix.
Table 1 presents the reduction in image resolution after applying our APS module, which demonstrates that APS significantly decreases the resolution of images. Some visual results are available in the Appendix. It is worth noting that APS is primarily designed for document images, where redundant regions are commonly distinctly defined. In contrast, elements in natural scene images tend to be more complex, making it challenging to identify extensive redundant regions. Consequently, APS generally has a limited impact on natural scene tasks. In the subsequent experimental section of our paper, we demonstrate that the application of APS consistently improves performance on multiple document image benchmarks without adversely affecting tasks involving natural scenes.
| Datasets | Adaptive Pixel Slimming | Avg. Resolution | Reduction |
| DocVQA | ✗ | 17482088 | - |
| ✓ | 15311524 | 36.0% | |
| InfoVQA | ✗ | 11812541 | - |
| ✓ | 11202193 | 18.1% |
3.2 More Flexible Image Encoding
@@
Most existing MLLMs [29] utilize a fixed rectangular resolution, such as 224224 or 336336. This approach distorts the original aspect ratio of the images, leading to text deformation and inefficiencies when upscaling images smaller than the preset resolution. While some methods [74, 76] allow for more flexible aspect ratios and resolutions, they often employ a patch-based processing strategy. Such a strategy may be suitable for natural scene images where the continuity and context of visual elements are less critical. However, it is less effective for document images [37] that contain dense text. Segmenting such images can disrupt the textual content, severing lines of text or splitting words and sentences across different patches. This discontinuity can impair the ability to accurately recognize and interpret textual information essential for effective document image understanding.
@@
We employ a more flexible approach that can accommodate varying resolutions and aspect ratios without the need for image patching. Specifically, for an image obtained by APS with the size of HW, we set a maximum allowable number of pixels, MAX_SIZE. If the product of H and W is less than MAX_SIZE, the image will not be resized. Otherwise, a scaling factor is calculated, and then the image is resized to rHrW. Here denotes the floor function to ensure that the resulting image does not exceed the maximum allowable pixels. One advantage of setting a maximum number of pixels rather than fixed dimensions is the ability to support more varied resolutions and preserve the original aspect ratio. Additionally, this manner avoids the loss of image details due to resolution reduction as much as possible without increasing computational resource consumption. For example, if we set the maximum pixels number MAX_SIZE as 960960, we can accommodate an input resolution of 1280720 without alteration. In contrast, a fixed-dimension strategy would require resizing the image to 960960, resulting in a 25% compression of the long side. Both strategies consume comparable computational resources since the number of visual tokens remains consistent.
@@
Although ViT pre-trained with CLIP [48] demonstrates impressive generalization capabilities for natural scene images, they perform suboptimally on dense document images [28]. Furthermore, their self-attention mechanism, which has a computational complexity that scales quadratically with image resolution, makes it challenging to support high-resolution inputs. Consequently, we opt to use the more efficient Swin Transformer [38] as our visual encoder, which has been pre-trained on full-text recognition tasks [22]. It enables us to process entire images as single inputs.
3.3 Dynamic Token Slimming
@@
Dual-center K-Means Clustering. To distinguish tokens into essential and nonessential types without introducing additional parameters, we apply K-Means clustering [39] with two centers to the output sequence from the visual encoder, as illustrated in Fig 1. However, after obtaining the two clusters, it is unclear which cluster contains the essential tokens and which contains the non-essential tokens. Drawing on findings from [37], nonessential tokens typically lack uniqueness and are often similar to other tokens. Therefore, we identify potential nonessential tokens by evaluating the similarity of each visual token with all other ones. The cluster with fewer of these nonessential tokens is determined to be the one containing the essential tokens. For more implementation details, please refer to Algorithm 1.
Similarity Weighted Aggregation. After obtaining essential and nonessential tokens, how to handle the nonessential tokens becomes a crucial issue. Inspired by [24, 30, 71], to mitigate potential clustering errors and prevent information loss, we do not discard the nonessential tokens outright. Instead, we aggregate them into the essential tokens to minimize the potential loss of effective information while enhancing the model’s processing efficiency. As illustrated in Fig. 3, for each nonessential token from the clustering step, we find the most similar essential token. Each essential token may correspond to zero or multiple nonessential tokens. If an essential token has corresponding nonessential tokens, we perform a similarity weighted summation of these tokens to obtain an aggregated token.

### v3

@@ Abstract
Current multimodal large language models (MLLMs) face significant challenges in visual document understanding (VDU) tasks due to the high resolution, dense text, and complex layouts typical of document images. These characteristics demand a high level of detail perception ability from MLLMs. While increasing input resolution improves detail perception capability, it also leads to longer sequences of visual tokens, increasing computational costs and straining the models’ ability to handle long contexts. To address these challenges, we introduce DocKylin, a document-centric MLLM that performs visual content slimming at both the pixel and token levels, thereby reducing token sequence length in VDU scenarios. We introduce an Adaptive Pixel Slimming (APS) preprocessing module to perform pixel-level slimming, increasing the proportion of informative pixels. Moreover, we propose a novel Dynamic Token Slimming (DTS) module to conduct token-level slimming, filtering essential tokens and removing others to adaptively create a more compact visual sequence. Experiments demonstrate DocKylin’s promising performance across various VDU benchmarks and the effectiveness of each component.
Introduction
@@
With the substantial growth in data and model sizes, as well as alignment with human preferences, current Large Language Models (LLMs) (OpenAI 2022; Touvron et al. 2023; The Vicuna Team 2023; Bai et al. 2023a; Team 2023; Yang et al. 2023) have demonstrated remarkable capabilities in reasoning, common sense understanding, expertise in specialized knowledge, zero/few-shot learning, and instruction following, shedding light on the potential for Artificial General Intelligence. Multimodal Large Language Models (MLLMs) aim to enhance these powerful LLMs by endowing them with multimodal capabilities, enabling the integration of diverse modalities beyond text alone.
@@
Although there has been significant progress in the development of MLLMs, their performance in visual document understanding (VDU) tasks remains suboptimal (Liu et al. 2023b; Zhang et al. 2024). A primary reason for this is their support for input resolutions of only up to , which is inadequate for high-resolution, fine-grained document images. Numerous efforts (Yu et al. 2024; Liu et al. 2024b; Li et al. 2023b; Wei et al. 2023a; Bai et al. 2023b; Ye et al. 2023a; Lin et al. 2023; Wei et al. 2024) have recently been made to increase the input resolution to address this limitation and achieved notable improvements. However, these efforts still face several challenges: 1) Increasing the resolution also escalates the redundant regions within documents, which results in an increased number of redundant visual tokens, complicating the task for LLMs to identify the correct answers. 2) Some methods (Li et al. 2023b; Wei et al. 2023a, 2024; Bai et al. 2023b) employ a fixed rectangular resolution, which can lead to text distortion given the typically large aspect ratios of document images. Moreover, scaling smaller images to larger ones results in inefficiencies. 3) To manage the extended sequences of visual tokens resulting from high resolution, most methods (Li et al. 2023b; Yu et al. 2024; Liu et al. 2024b; Bai et al. 2023b) employ a fixed compression ratio or extract a predetermined number of tokens. While this approach may be suitable for natural scene images, it is less effective for document scenarios where content density varies significantly. For example, images of scientific papers are information-dense, whereas images of identification cards are comparatively sparse.
@@
To tackle these challenges, we propose DocKylin, a document-centric MLLM that incorporates several innovative designs: Firstly, we propose a parameter-free Adaptive Pixel Slimming (APS) preprocessing technique that utilizes gradient information to identify and eliminate redundant regions within document images, thereby reducing the proportion of redundant pixels and enhancing computational efficiency. Secondly, we adopt a more flexible image encoding strategy by capping the maximum number of pixels instead of fixing resolutions. This allows for variable resolutions and aspect ratios. Our lightweight Swin encoder (Liu et al. 2021) processes high-resolution document images holistically, ensuring feature coherence. Thirdly, we propose a Dynamic Token Slimming (DTS) method based on a dual-center clustering approach, which is also parameter-free. DTS filters useful tokens from a vast pool of visual tokens, obtaining visual sequences with reduced and dynamic lengths. Incorporating these innovations, we developed our DocKylin, which demonstrates superior performance across multiple VDU benchmarks. Extensive experiments demonstrate that DocKylin outperforms existing methods across various VDU benchmarks. Furthermore, both the proposed APS and DTS are parameter-free and modular, facilitating easy integration into existing MLLMs to achieve enhanced performances.
Related Works
Visual Document Understanding
Visual Document Understanding (VDU) (Cui et al. 2021; Zhang et al. 2023; Xu et al. 2020; Kim et al. 2022; Li et al. 2021) centers on the automated processing, interpretation, classification, and extraction of information from document images with varied typesetting formats. This task is particularly challenging due to the diverse layouts, often poor image quality, and complex structures within the documents.
Current advanced VDU methods can be broadly categorized into OCR-dependent and OCR-free approaches. OCR-dependent methods (Huang et al. 2022; Xu et al. 2021, 2020; Appalaraju et al. 2021, 2024; Tang et al. 2023; Gu et al. 2022; Wang, Jin, and Ding 2022; Li et al. 2021) leverage external Optical Character Recognition engines to extract text and layout information from documents. Such additional OCR step complicates the workflow and makes performance heavily reliant on OCR accuracy. Errors in OCR step can be accumulated to the subsequent understanding model. In contrast, OCR-free methods (Kim et al. 2022; Davis et al. 2022; Yu et al. ) process document images directly as input and produce end-to-end results without the need for an additional OCR step.
Text-centric MLLMs
@@
Leveraging large language models (LLMs) trained on extensive datasets, emerging multimodal LLMs have recently demonstrated significant potential in visual document understanding (VDU) tasks. However, their performance still falls short of expectations (Liu et al. 2023b; Shi et al. 2023; Tang et al. 2024b). Therefore, there has been a surge of recent work studying how to enhance the perception, recognition, and understanding abilities of MLLMs in such challenging scenario. Some approaches (Tanaka et al. 2024; Kim et al. 2023) leverage additional Optical Character Recognition (OCR) engines to recognize text in images and input it into MLLMs. Others (Li et al. 2023b; Liu et al. 2024b; Bai et al. 2023b; Dong et al. 2024; Feng et al. 2023a; Ye et al. 2023a; Wei et al. 2023a) enhance the fine-grained perceptual capabilities of MLLMs by maintaining the high resolution of the input images. This includes strategies such as splitting the image into small pieces to accommodate the VIT’s requirement for fixed and small-size inputs and utilizing visual encoders that are more suitable for text images. Authors of (Feng et al. 2023b; Wang et al. 2023) explore learning text grounding for MLLMs, while (Hu et al. 2024; Tang et al. 2024a) investigate the use of more superior and larger scale instruction data to boost performance. Although these efforts have significantly improved the performance of MLLMs on text-centric tasks, there is still substantial room for improvement.
Visual Token Compression in MLLMs
@@
Maintaining high-resolution input images is a straightforward and effective method to enhance the performance of MLLMs on text images. However, increased resolutions inevitably lead to a longer visual sequence. There are various methods proposed for compressing the length of visual sequences in current MLLMs. One of the earlier and commonly used approaches (Dai et al. 2024; Li et al. 2023a; Alayrac et al. 2022) involves utilizing a group of learnable query tokens to extract information in a cross-attention manner. Another method (Chen et al. 2023a; Yu et al. 2024) involves concatenating adjacent tokens along the channel dimension to reduce sequence length. Additionally, convolutional neural networks (CNNs) have also been proven effective for visual token compression (Cha et al. 2023; Hu et al. 2024). However, all these methods either extract tokens of a fixed length or compress them at a fixed ratio, which is unsuited for document images, where the content density can vary significantly. A token sequence that is too short or a compression ratio that is too high can result in an inadequate representation of some document images. Conversely, longer token sequences or lower compression rates can reduce computational efficiency.
@@
Recent studies (Cai et al. 2024) have shown that the optimal length of visual token sequences varies for each sample. Consequently, some recent approaches have explored the adaptive compression of visual sequences. Most of these methods (Liu et al. 2024a; Shang et al. 2024; Chen et al. 2024; Lin et al. 2024) dynamically discard nonessential tokens using attention scores between visual tokens or between visual tokens and class tokens. However, considering the differences between text and natural objects with complete semantics, this paradigm may not be suitable for text images. For text-centric images, HRVDA (Wei et al. 2024) employs text position annotations to train a separate visual token filtering network, which requires introducing additional parameters and training procedures. Compared to the methods above, we explore alternative dynamic compression schemes specifically for document images, which include pixel-level redundancy removal and unsupervised clustering for filtering out nonessential tokens.
Methodology
@@
As illustrated in Fig. 1, the architecture of DocKylin incorporates an image encoder, an MLP layer that aligns the visual modality to the language modality, and a LLM. Building on this framework, DocKylin introduces the following enhancements: 1) An Adaptive Pixel Slimming (APS) module preprocesses images before they are fed into the image encoder, which can removes redundant regions and reduces resolution while increasing the proportion of informative pixels. 2) a more lightweight and flexible image encoder by capping the maximum number of pixels instead of fixing resolutions. 3) A Dynamic Token Slimming (DTS) module handles the visual tokens produced by the visual encoder and MLP layer. DTS further removes nonessential tokens to enhance both performance and efficiency.
Adaptive Pixel Slimming
@@
Perceiving textual information in images typically requires maintaining a high resolution of input, especially for text-rich document images. This poses significant challenges for current resource-intensive MLLMs. To address this issue, existing methods employ strategies such as dividing the images into patches for processing (Liu et al. 2024b; Ye et al. 2023a), employing visual encoders that support high resolution (Wei et al. 2023a), leveraging OCR results as additional inputs (Tanaka et al. 2024; Kim et al. 2023), or applying compression transformations to the images beforehand (Feng et al. 2023a).
@@
Here, we propose a novel approach called Adaptive Pixel Slimming (APS) to efficiently reduce the resolution of document images, which is orthogonal to the approaches mentioned above. APS is based on the observation that, as illustrated in Fig. 2, while document images usually exhibit high resolution, they commonly contain a significant proportion of redundant regions. These regions not only increase the computational burden on the visual encoder but may also pose challenges (Liu et al. 2024b; Wei et al. 2024) to the reasoning capabilities of the LLMs by the increased redundant visual tokens. APS is thus designed to identify and remove these regions. Specifically, as shown in Fig. 2, during the gradient extraction stage, we utilize the Sobel operator to extract the gradient map from the original image . Regions with low gradient values are considered smooth and devoid of significant information. Therefore, in the redundant region determination stage, we identify the redundant regions in the horizontal and vertical directions by locating areas with consistently low gradient values. For instance, to identify redundant regions in the image’s horizontal direction: if the sum of gradient values for a specific row is below a predefined threshold, we classify that row as redundant. A sequence of contiguous redundant rows constitutes a redundant region in the horizontal direction. The computation for the vertical direction follows a similar procedure. The final image can be obtained by removing the identified redundant rows and columns. For more details on the algorithm and parameter settings, please refer to the supplementary materials.
Flexible and Lightweight Image Encoder
@@
Most existing MLLMs (Li et al. 2023b) utilize a fixed rectangular resolution, such as 224224 or 336336. This approach distorts the original aspect ratio of the images, leading to text deformation and inefficiencies when upscaling images smaller than the preset resolution. While some methods (Ye et al. 2023a; Yu et al. 2024) allow for more flexible aspect ratios and resolutions, they often employ a patch-based processing strategy. Such a strategy may be suitable for natural scene images where the continuity and context of visual elements are less critical. However, it is less effective for document images (Liu et al. 2024b) that contain dense text. Segmenting such images can disrupt the textual content, severing lines of text or splitting words and sentences across different patches. This discontinuity can impair the ability to accurately recognize and interpret textual information essential for effective document image understanding.
@@
We employ a more flexible approach that can accommodate varying resolutions and aspect ratios without the need for image patching. Specifically, for an image obtained by APS with the size of , we set a maximum allowable number of pixels, MAX_SIZE. If the product of H and W is less than MAX_SIZE, the image will not be resized. Otherwise, a scaling factor is calculated, and then the image is resized to . Here denotes rounding down to the nearest integer to ensure that the resulting image does not exceed the maximum allowable pixels. One advantage of setting a maximum number of pixels rather than fixed dimensions is the ability to support more varied resolutions and preserve the original aspect ratio. Additionally, this manner avoids the loss of image details due to resolution reduction as much as possible without increasing computational resource consumption. For example, if we set the maximum pixel number MAX_SIZE as 960960, we can accommodate an input resolution of 1280720 without alteration. In contrast, a fixed-dimension strategy would require resizing the image to 960960, resulting in a 25% compression of the long side. Both strategies consume comparable computational resources since the number of visual tokens remains consistent.
@@
Although ViT pre-trained with CLIP (Radford et al. 2021) demonstrates impressive generalization capabilities for natural scene images, they perform suboptimally on dense document images (Li et al. 2024). Furthermore, their self-attention mechanism, which has a computational complexity that scales quadratically with image resolution, makes it challenging to support high-resolution inputs. Consequently, we opt to use the more efficient Swin Transformer (Liu et al. 2021) as our visual encoder, which has been pre-trained on full-text recognition tasks (Kim et al. 2022). It enables us to process an entire image as a single input.
Dynamic Token Slimming
@@
Dual-center K-Means Clustering. To distinguish tokens into essential and nonessential types without introducing additional parameters, we apply K-Means clustering (Lloyd 1982) with two centers to the output sequence from the visual encoder, as illustrated in Fig 1. However, after obtaining the two clusters, it is unclear which cluster contains the essential tokens and which contains the non-essential tokens. Drawing on findings from (Liu et al. 2024b), nonessential tokens typically lack uniqueness and are often similar to other tokens. Therefore, we identify potential nonessential tokens by evaluating the similarity of each visual token with all other ones. The cluster with fewer of these nonessential tokens is determined to be the one containing the essential tokens. For more implementation details, please refer to Algorithm 1.
| Methods | Vis. Encoder (Params.) | Decoder | DocVQAval | InfoVQAval | ChartQA | FUNSD | SROIE | POIE |
| mPLUG-Owl (Ye et al. 2023b) | CLIP-L (0.3B) | LLaMA-7B | 7.4 | 20.0 | 7.9 | 0.2 | 0.1 | 0.3 |
| InstructBLIP (Dai et al. 2024) | CLIP-G (2B) | Vicuna-7B | 4.5 | 16.4 | 5.3 | 0.2 | 0.6 | 1.0 |
| LLaVA1.5 (Liu et al. 2023a) | CLIP-L (0.3B) | LLaMA-7B | 8.5 | 14.7 | 9.3 | 0.2 | 1.7 | 2.5 |
| Qwen-VL (Bai et al. 2023b) | CLIP-G (2B) | Qwen-7B | 48.1 | 23.9 | 53.4 | 23.9 | 34.5 | 20.6 |
| Monkey (Li et al. 2023b) | CLIP-G (2B) | Qwen-7B | 50.1 | 25.8 | 54.0 | 24.1 | 41.9 | 19.9 |
| InternVL (Chen et al. 2023b) | IViT (6B) | Vicuna-7B | 28.7 | 23.6 | 45.6 | 6.5 | 26.4 | 25.9 |
| InternLM-XComposer2 (Dong et al. 2024) | CLIP-L (0.3B) | InernLM2-7B | 39.7 | 28.6 | 51.6 | 15.3 | 34.2 | 49.3 |
| DocKylin (ours) | Swin (0.07B) | Qwen-7B | 65.1 | 34.8 | 59.0 | 25.5 | 49.5 | 36.1 |
| Methods | Visual Encoder (Params.) | Decoder | DocVQAtest | InfoVQAtest | ChartQA | VisualMRC | DeepForm | WTQ |
| Donut (Kim et al. 2022) | Swin (0.07B) | BERT(1-4 layers) | 67.5 | 11.6 | 41.8 | 93.9 | 61.6 | 18.8 |
| Pix2Struct (Lee et al. 2023) | - | Transformers-1.3B | 76.6 | 40.0 | 58.6 | - | - | - |
| mPLUG-Doc (Ye et al. 2023b) | CLIP-L (0.3B) | LLaMA-7B | 62.2 | 38.2 | 57.4 | 188.8 | 42.6 | 26.9 |
| UReader (Ye et al. 2023a) | CLIP-L (0.3B) | LLaMA-7B | 65.4 | 42.2 | 65.7 | 221.7 | 49.5 | 29.4 |
| DocPeida (Feng et al. 2023a) | Swin (0.1B) | Vicuna-7B | 47.1 | 15.2 | 46.9 | - | - | - |
| HRVDA (Wei et al. 2024) | Swin | LLaMA2-7B | 72.1 | 43.5 | 67.6 | 211.5 | 63.2 | 31.2 |
| Vary (Wei et al. 2023a) | ViTDet-B & CLIP-L (0.4B) | Vicuna-7B | 76.3 | - | 66.1 | - | - | - |
| TextMonkey (Liu et al. 2024b) | CLIP-G (2B) | Qwen-7B | 71.5 | - | 65.5 | - | 61.6 | 30.6 |
| TextHawk (Yu et al. 2024) | SigLIP-SO (0.4B) | InternLM-7B | 76.4 | 50.6 | 66.6 | - | - | 34.7 |
| LayoutLLM (Huang et al. 2022) | LayoutLMv3-L (0.3B) | LLaMA2-7B | 74.3 | - | - | - | - | - |
| DocKylin (ours) | Swin (0.07B) | Qwen-7B | 77.3 | 46.6 | 66.8 | 319.9 | 64.2 | 32.4 |
Similarity Weighted Aggregation. After obtaining essential and nonessential tokens, how to handle the nonessential tokens becomes a crucial issue. Inspired by (Kong et al. 2022; Liang et al. 2021; Wei et al. 2023b), to mitigate potential clustering errors and prevent information loss, we do not discard the nonessential tokens outright. Instead, we aggregate them into the essential tokens to minimize the potential loss of effective information while enhancing the model’s processing efficiency. As illustrated in Fig. 3, for each nonessential token from the clustering step, we find the most similar essential token. Each essential token may correspond to zero or multiple nonessential tokens. If an essential token has corresponding nonessential tokens, we perform a similarity weighted summation of these tokens to obtain an aggregated token.